If the average weight of fish caught for the day is 36 pounds with a 4 point deviation, what is the probabilty that if 30 fishermen were checked the average weight would be less than 31.5?
1
Expert's answer
2011-07-21T05:43:02-0400
As the deivation is known, we use one-tailed z-test:
z = (X - μ)/(σ / √ n) = (31.5 - 36)/ (4/ √30) = 6.02

The probability is approximately 0.